# HopCast: Calibration of Autoregressive Dynamics Models

### v4

@@ A.3 Moment Matching
where $\alpha=\frac{1-z}{2}$ for $z\%$ prediction intervals. The $\Phi^{-1}$ denotes the inverse cumulative distribution function of the standard normal distribution. In contrast, equations 3 & 4 were used to generate prediction intervals with HopCast.
@@ Appendix B Dynamical Systems
B.1 Lotka-Volterra (Wangersky, 1978)
@@
B.3 FitzHugh-Nagumo (FHN) (Izhikevich & FitzHugh, 2006)
@@ B.4 Lorenz95 (Lorenz, 1996)
B.5 Glycolytic Oscillator (Daniels & Nemenman, 2015)
@@ Appendix C Hyperparameters
In Table 3, we provide the Sequence Length ($\mathtt{S_{L}}$) for each output of the Predictor and a number of models in Probabilistic Ensembles for each experimental setting in Table  2.
@@
Appendix D Ablation Studies
D.1 Encoder Architecture
@@
D.2 Initial condition and context state ($c_{t}^{n}$)
As discussed in section 1, our context state ($c_{t}^{n}$) comprises the initial condition, the current state of the system, and the timestep ID. Here, we show the results on two systems without the initial condition a part of the context state in Table  5. The MSE increases significantly for both cases, while the CE exhibits minor changes. This shows the significance of making the initial condition a part of the context state.
@@ D.2 Initial condition and context state ($c_{t}^{n}$)
D.3 Size of Association Memory
@@ D.4 Embedding dimension
D.5 Impact of shuffled data
@@
D.6 Number of models in Probabilistic Ensembles
@@
D.7 Sequence length ($\mathtt{S_{L}}$) & Calibration
@@
Appendix E Implementation Details
@@
E.1 Baselines Implementation
We train a population of 15 models for the Probabilistic Ensembles. Regarding architecture, two layers of a fully connected feedforward model with 400 neurons each were used for LV and FHN, and three layers with 400 neurons for Lorenz, Lorenz95, and Glycolytic Oscillator. The train/test datasets were prepared following an 80/20 split. The batch size, learning rate, optimizer, and epochs were kept the same for all experiments, and are 128, 0.001, $\mathtt{Adam}$ (Kingma & Ba, 2017), and 1000, respectively. The early stopping was used as a criterion to train the Probabilistic Ensembles (Yao et al., 2007). The number of particles $P$ for each model in Probabilistic Ensembles was 1 for Expectation, and 20 for Moment Matching & Trajectory Sampling (Chua et al., 2018).
E.2 HopCast Implementation
The Encoder was a fully connected feedforward model with one layer and 100 neurons. The section 7 contains details about tuning an important hyperparameter, i.e., $\mathtt{S_{L}}$, of HopCast. Table 3 has $\mathtt{S_{L}}$ for each output of all systems at various noise scaling factors $\sigma$. The learning rate of 0.001 and the optimizer $\mathtt{AdamW}$ (Loshchilov & Hutter, 2019) were kept the same for all experiments. The batch size and epochs were different for each experiment, and are provided in the form of $\mathtt{yml}$ files as a supplementary material along with other hyperparameters for each system and noise scaling factor ($\sigma$).

### v5

@@ A.3 Moment Matching
where $\alpha=\frac{1-z}{2}$ for $z\%$ prediction intervals. The $\Phi^{-1}$ denotes the inverse cumulative distribution function of the Standard Normal distribution. In contrast, equations 5 & 6 were used to generate prediction intervals with HopCast.
@@ Appendix B Dynamical Systems
B.1 Lotka-Volterra (wangersky1978lotka)
@@ B.2 Lorenz (brunton2016discovering)
B.3 FitzHugh-Nagumo (FHN) (Izhikevich:2006)
@@ B.4 Lorenz95 (lorenz1996predictability)
B.5 Glycolytic Oscillator (daniels2015efficient)
@@ Appendix C Hyperparameters
In Table 3, we provide the Sequence Length ($\mathtt{S_{L}}$) for each output of the Predictor and a number of models in Probabilistic Ensembles for each experimental setting in Table  1.
@@
Appendix D Data Generation
Table 4: Details about data generation
Model $\Delta t$ Timesteps Trajectories
Lotka Volterra 0.1 300 500
Lorenz 0.01 300 1000
FHN 0.5 400 350
Lorenz95 0.01 300 666
Glycolytic 0.01 400 750
Appendix E Ablation Studies
E.1 Encoder Architecture
@@
E.2 Initial condition and context state ($c_{t}^{n}$)
As discussed in section 1, our context state ($c_{t}^{n}$) comprises the initial condition, the current state of the system, and the timestep ID. Here, we show the results on two systems without the initial condition a part of the context state in Table  6. The MSE increases significantly for both cases, while the CE exhibits minor changes. This shows the significance of making the initial condition a part of the context state.
@@ E.2 Initial condition and context state ($c_{t}^{n}$)
E.3 Number of Retrieved Samples ($s$)
To generate calibrated prediction intervals, a set of $s$ samples is drawn from $\mathbf{a}_{x}^{t}$ as described in section 4.3. Here, we analyze the sensitivity of proposed metrics with varying values of $s=\{100,500,800,1000\}$ for two settings from Table 1. The results everywhere else in the paper are reported with $s=1000$. In Table 7, it can be seen that the results are largely unchanged as the number of samples varies from 1000 to 500. However, at $s=100$, we see a drop in performance in terms of MSE and CE, suggesting that too few samples limit the diversity of error samples, negatively impacting the performance.
Table 7: Comparison of proposed metrics with varying number of retrieved residual samples ($s$) from MHN memory. The results are averaged over three runs.
Dynamical System Glycolytic Oscillator Lotka Volterra Noise Level $\bm{\sigma=0.1}$ $\bm{\sigma=0.1}$ Metrics MSE PI-Width CE MSE PI-Width CE 1000 $0.072\pm 0.003$ $25\pm 0.007 0.0$$018\pm 0.006$ $6.10\pm 0.20$ $1.88\pm 0.05$ $0.0070\pm 0.002$ 800 $0.070\pm 0.005$ $24\pm 0.009 0.0$$018\pm 0.008$ $6.10\pm 0.2$ $1.89\pm 0.03$ $0.007\pm 0.003$ 500 $0.071\pm 0.005$ $26\pm 0.003 0.0$$015\pm 0.008$ $6.13\pm 0.20$ $1.87\pm 0.04$ $0.006\pm 0.005$ 100 $0.096\pm 0.005$ $18\pm 0.006 0.0$$025\pm 0.008$ $7.9\pm 0.17$ $69\pm 0.04 0.0$$019\pm 0.004$
E.4 Top-$k$ retrieval
As discussed in section 4.3, the $s$ samples are drawn from $\mathbf{a}_{x}^{t}$ to generate calibrated prediction intervals. Here, we analyze an alternative approach to sampling, i.e., using top-$k$ probabilities from $\mathbf{a}_{x}^{t}$. Table 8 shows the results on proposed metrics with varying values of $k=\{100,500,800,1000\}$. The results remain largely unchanged for $k\geq 500$. At $s=100$, the CE shows an increase for both settings, indicating that too few samples lead to decreased diversity in sampled errors.
Table 8: Comparison of proposed metrics based on top-$k$ probabilities from $\mathbf{a}_{t}^{x}$ instead of sampling. The results are reported for $k=\{100,500,800,1000\}$. The results are averaged over three runs.
Dynamical System Glycolytic Oscillator Lotka Volterra Noise Level $\bm{\sigma=0.1}$ $\bm{\sigma=0.1}$ Metrics MSE PI-Width CE MSE PI-Width CE 1000 $0.67\pm 0.002$ $24\pm 0.007 0.0$$011\pm 0.005$ $7.31\pm 0.99$ $2.38\pm 0.33$ $0.016\pm 0.003$ 800 $0.067\pm 0.002$ $24\pm 0.01 0.0$$011\pm 0.005$ $7.52\pm 1.18$ $2.41\pm 0.38$ $0.014\pm 0.004$ 500 $0.069\pm 0.003$ $24\pm 0.009 0.0$$017\pm 0.005$ $8.39\pm 2.50$ $2.53\pm 0.64$ $0.017\pm 0.008$ 100 $0.07\pm 0.004$ $20\pm 0.005 0.0$$06\pm 0.03$ $6.36\pm 0.22$ $1.24\pm 0.11$ $0.25\pm 0.036$
E.5 Size of Association Memory
@@ E.6 Embedding dimension
E.7 Impact of shuffled data
@@
E.8 Number of models in Probabilistic Ensembles
@@
E.9 Sequence length ($\mathtt{S_{L}}$) & Calibration
@@
Appendix F Implementation Details
@@
F.1 Baselines Implementation
We train a population of 15 models for the Probabilistic Ensembles. Regarding architecture, two layers of a fully connected feedforward model with 400 neurons each were used for LV and FHN, and three layers with 400 neurons for Lorenz, Lorenz95, and Glycolytic Oscillator. The train/test datasets were prepared following an 80/20 split. The batch size, learning rate, optimizer, and epochs were kept the same for all experiments, and are 128, 0.001, $\mathtt{Adam}$ (kingma2017adammethodstochasticoptimization), and 1000, respectively. The early stopping was used as a criterion to train the Probabilistic Ensembles (yao2007early). The number of particles $P$ for each model in Probabilistic Ensembles was 1 for Expectation, and 20 for Moment Matching & Trajectory Sampling (chua2018deep).
F.2 HopCast Implementation
The Encoder was a fully connected feedforward model with one layer and 100 neurons. The section 7 contains details about tuning an important hyperparameter, i.e., $\mathtt{S_{L}}$, of HopCast. Table 3 has $\mathtt{S_{L}}$ for each output of all systems at various noise scaling factors $\sigma$. The learning rate of 0.001 and the optimizer $\mathtt{AdamW}$ (loshchilov2019decoupledweightdecayregularization) were kept the same for all experiments. The batch size and epochs were different for each experiment, and are provided in the form of $\mathtt{yml}$ files as a supplementary material along with other hyperparameters for each system and noise scaling factor ($\sigma$).